OpenStudy (anonymous):

A sealed rectangular container 6 cm by 12 cm by 15 cm is sitting on its smallest face. It is filled with water up to 5 cm from the top. How many centimeters from the bottom will the water level reach if the container is placed on its largest face?

OpenStudy (anonymous):

the totl volums of the water when it is sitting on the smallest face would be 6*5*10 (since the height from the bottom would be 15-5) = 300 cubic cm Now when the container is placed on its largest face, the volume of water is still is the same i.e 300. the equation becomes 15*12*h=300 h=20/12 i.e 5/3 = 1.66 from the bottom or (6-1.67)cm from the top

OpenStudy (anonymous):

i don't get it